what is the equation of the line going through (5,3) and (7,-8)

Found 2 solutions by checkley71, tutorcecilia:
Answer by checkley71(8403)

SLOPE OF THIS LINE IS (-8-3)/(7-5) OR -11/2 OR -5.5
Y INTERCEPT IS 3=-5.5*5+B OR 3=-27.5+B OR B=30.5
THEN Y=-5.5X+30.5

Answer by tutorcecilia(2152)

(5,3) and (7,-8) [Points on a line]
m = (y2-y1)/x2-x1) [Formula for the slope of a line]
m = ((-8)-3)/(7-5) [Plug-in the values]
m = -11/2 [The slope of the line]
.
y = mx+b [The slope-intercept formula for the equation of a line]
y = the y-value of a point on the line
m = the slope of the line
x = the corresponding x-value for a point on the line
b = the y-intercept; it is where the line crosses the y-axis (0, b)
.
y = mx + b
3 = (-11/2)(5) + b [Pick a point and plug-in the values]
3 = (-55/2) + b [Solve for "b"]
(55/2) + (3) = b
61/2 = b
.
y=(-11/2)x + (61/2) [Plug-in the values of the slope and y-intercept][Final answer]
.
Check by plugging the values for either of the original points